The company founder hires us as consultants and asks that we oversee the accounting for new equipment purchased on January 1.Actual & Estimated Units-of-Production Year 1 Production Actual 35.000 Units Year 2 Production Estimated 55.000 Units Year 3#tableau € → ſ oņO 1(a). Determine the equipments first-year depreciation under the straight-line method. 1(b). Determine th#tableau E → K of TO 1(a). Determine the equipments first-year depreciation under the straight-line method. (b). Determine ttableau * » K ę po 1(a). Determine the equipments first-year depreciation under the straight-line method. 1(b). Determine th1(a). Determine the equipments first-year depreciation under the straight-line method. 1(b). Determine the equipments first1(a). Determine the equipments first-year depreciation under the straight-line method. 1(b). Determine the equipments first1(a). Determine the equipments first-year depreciation under the straight-line method. 1(b). Determine the equipments first
Answer #1

Part 1 A

Straight-line method

Choose numerator

/

Choose denominator

=

Annual depreciation expense

Cost minus salvage

/

Estimated useful life (years)

=

Depreciation expense

$30000

/

4

=

$7500

Cost minus salvage = depreciable base = 40000-10000 = 30000

Part 1 B

Units-of-production depreciation

Choose numerator

/

Choose denominator

=

Annual depreciation expense

Cost minus salvage

/

Total estimate units of production

Depreciation expense per unit

$30000

/

120000

$0.40

Year

Annual production (units)

Deprecation expense

1

35000

$14000

Total estimated production = 35000+55000+25000+5000 = 120000

Part 1 C

Depreciation for the period

End of period

Annual period

Beginning of period book value

Deprecation rate (%)

Depreciation expense

Accumulated depreciation

Book value

First year

40000

50%

20000

20000

20000

Deprecation rate = 1/ estimated life * 2 = ¼ *2 = 50%

End period book value = beginning book value – depreciation expense

Part 2

Which method in part 1 results in the highest net income in the first year

Straight-line method

The method with the lowest depreciation expense will give the highest net income.

Part 3

Which method would we recommend the company use?

Units of production method

Because according this method, the equipment will be depreciated in the proportion of units produced. It will help to derive fair amount of depreciation.

Part 4

Which method would result in the highest amount of depreciation over an asset’s useful life?

Double-declining balance method

Because it ignores the salvage value.